A) Derive the following relationship for interpolation U= =U+U2 (6 Marks)

pls
Answer #1

A < > Sense В 2,0) A (انا رومه PL We know that equalion of live (y- y) = n(x-1) From the plot m = Hope Yo-Y (U-U)= m (P-P.) UFormula for interpolation has been evaluated.